OpenStudy (anonymous):

compute the difference quotient of the given function g(t)=42

OpenStudy (anonymous):

difference quotient at the given point \(x\) equals \[\frac{g(x+\Delta x)-g(x)}{\Delta x}\]

OpenStudy (anonymous):

but \(g(t)\) is a constant function can u tell me what will be the value of \(g(t+\Delta t)\) or \(g(t)\)

OpenStudy (anonymous):

42

OpenStudy (anonymous):

thats right so what will be the difference quotient

OpenStudy (anonymous):

x

OpenStudy (anonymous):

wait.... i don't understand you're question

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

well we know that \( g(\text{anything})=42\)

OpenStudy (anonymous):

difference quotient \[\frac{42-42}{\Delta t}\]

OpenStudy (anonymous):

my equation is f(x+h)-f(x)/h

OpenStudy (anonymous):

oh sorry i mean \(h\) by \(\Delta x\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would be Q(42+h)-Q(42)/h ?

OpenStudy (anonymous):

\(x=42\) ? or \(g(x)=42\)

OpenStudy (anonymous):

u have \[\frac{g(x+h)-g(x)}{h}=\frac{42-42}{h}=0\]since \(g\) is a constant function

OpenStudy (anonymous):

so difference quotient is 0

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

would H(x)=2z^2+9 be solved the same way?

OpenStudy (anonymous):

@mukuchla